### v1

[1]organization=X Computational Physics Division, Los Alamos National Laboratory, city=Los Alamos, NM, postcode=87545, country=USA
[2]organization=Theoretical Division, Los Alamos National Laboratory, city=Los Alamos, NM, postcode=87545, country=USA
Phase field dislocation dynamics study of grain boundary-dislocation interactions.
Summary
A new phase field dislocation dynamics formulation is presented, which couples micromechanical solvers and the time-dependent Ginzburg-Landau equation. Grain boundary (GB)-dislocation interactions are studied by describing GBs as inclusions. Grain boundary properties are computed from Molecular Statics simulations and an additional contribution to the total energy that takes into account the GB energy is considered in the calculations. Interaction of a screw dislocation with minimum energy and metastable states of low and high angle 110 symmetric tilt grain boundaries are studied. We show good agreement between predictions from our phase field dislocation dynamics formulation and molecular dynamics simulations of grain boundary-dislocation interactions.
Keywords: Phase field method, Dislocation dynamics, Molecular Dynamics, Grain boundary, Slip Transfer.
1 Introduction
††LA-UR-25-26159New manufacturing processes such as Cold Spray [Murgas et al.,, 2024], Accumulative Roll Bonding (Severe Plastic Deformation) [Zhang et al.,, 2023], and Laser Powder Bed Fusion [Gray III et al.,, 2017] produce microstructures with complex and dense grain boundary (GB) networks, resulting in material response that varies considerably when compared to microstructures fabricated with classical manufacturing methods. During deformation, dislocations nucleate, annihilate and move through the grains encountering obstacles that affect their motion. As GBs are inherent in polycrystalline materials, they are a common impediment to dislocation motion and play an important role in controlling the material properties. The Hall-Petch relationship, which predicts that the flow stress is inversely proportional to the square root of the grain size, is the most common link between grain size (or GB density) and a mechanical property [Hall,, 1951, Petch,, 1953]. In other words, microstructures with small grains have more GBs that impede dislocation motion resulting in an increase of the yield stress. Conversely, microstructures with larger grains have lower yield stress.
However, experimental observations [Nieto-Valeiras et al.,, 2024] and molecular dynamics (MD) simulations show that certain GBs do not behave as strong obstacles to dislcocation motion [Dang et al.,, 2024]. GB-dislocation interactions vary, and dislocations can be absorbed, pinned, fully or partially transmitted when impinging on a GB [Lim and Raj,, 1985, Sutton,, 1995, Zaefferer et al.,, 2003, Priester,, 2012]. The type of interaction depends on the type of dislocation, applied thermomechanical load and nature of the GB, i.e., GB misorientation, GB structure, and GB plane. Hence, by predicting the GB-dislocation interaction and engineering ‘favorable’ GBs, it is possible to enhance and tailor the in-service properties of the materials, especially during the application of new manufacturing processes where GB networks are dense [Gray III et al.,, 2017]. Current criteria for dislocation transmission across GBs are based on experimental observations and consider geometric parameters, stress state and energetics, or combinations of these, in their functional forms. Geometric criteria are based on the alignment of the slip direction, plane and intersection line between the slip-plane normal and GB plane between incoming and outgoing slip systems. Examples of geometric criteria include: [Livingston and Chalmers,, 1957], [Shen et al.,, 1986], [Luster and Morris,, 1995], residual Burgers vector [Marcinkowski and Tseng,, 1970, Bollmann,, 2012], misorientation or disorientation [Aust and Chen,, 1954, Clark and Chalmers,, 1954] and [Werner and Prantl,, 1990]. Stress-based criteria consider the Schmid factor of the incoming and outgoing slip systems [Reid,, 2016], the generalized Schimd factor [Bieler et al.,, 2014] or the resolved shear stress of the incoming and outgoing slip systems [Lee et al.,, 1989]. Combinations of geometric parameters with the accumulated shear or the Schmid factor [Bieler et al.,, 2014] have been proposed. Lastly, other criteria consider the energy change of the interaction using a line tension model [Koning et al.,, 2002]. In this model, the energy of the incoming, outgoing and residual dislocations is consider, as well as GB dislocations and GB energy barriers [Tsuru et al.,, 2009, Sangid et al.,, 2011, Li et al.,, 2020, 2022].
Lower scale simulations have provided new insights into GB-dislocations interactions. MD simulations are ideal to study unit processes dictating GB-dislocation interactions over a wide range of GB structures and types of dislocation [Spearot and Sangid,, 2014, Adams et al.,, 2019, Dang et al.,, 2024, 2025, Suresh et al.,, 2023]. In Spearot and Sangid, [2014] and Adams et al., [2019], geometric criteria, residual Burgers vector and GB disorientation angle have been identified as indicators of dislocation transmission as observed experimentally; however, these studies were limited to minimum energy GB configurations. In Dang et al., [2024, 2025], GB-interactions were studied for stable and metastable GB configurations, with different reactions for stable and metastable GB configurations. These results indicate that GB structure, including minimum energy and metastable states, should be considered in the development of multiscale material models.
Due to time and length scale limitations of atomistic simulations, mesoscale models are necessary to upscale information and new mechanisms to the continuum scale. Examples of mesoscale models are Discrete Dislocation Dynamics (DDD) [Li et al.,, 2009], Coupled Atomistic/Discrete Dislocation (CADD) [Dewald and Curtin,, 2006, 2007, 2011, Tadmor et al.,, 1996, Shilkrot et al.,, 2002], and Concurrent Atomistic-Continuum (CAC) [Xu et al.,, 2016, Su et al.,, 2023, Peng et al.,, 2022, Chen and Lee,, 2005]. In DDD, dislocations are represented by segments and the precision of simulations highly depends on the rules defined a priori [Cai and Bulatov,, 2004, Bertin et al.,, 2024]. In Li et al., [2009], GBs were described as geometric interfaces and the interaction was simplified to permeable or impermeable. In Bamney et al., [2021], the authors presented a DDD framework in which the structure of stable and metastable GBs were represented using dislocation arrays. However, a transmission rule needs to be defined and, e.g., in Bamney et al., [2021], the authors used the LRB criterion to model transmission [Shen et al.,, 1986]. Moreover, DDD does not represent the stacking fault of dislocations which can impact the GB-dislocation interaction, i.e., partial transmission and extended stacking faults [Dang et al.,, 2024]. CADD depends on the definition of defects in the continuum domain and the treatment of the internal boundary force that connects the quasi-continuum and continuum defect problems without the introduction of ghost forces. CADD simulations with GBs have been limited to 2D studies, hence limiting their application to special GBs with specific symmetries [Dewald and Curtin,, 2006, 2007, 2011]. CAC has been applied to study GB-dislocaiton interactions in FCC metals [Xu et al.,, 2016, Su et al.,, 2023] and dislocation pileup in a bicrystal system with square and hexagonal lattice [Peng et al.,, 2022]. CAC simulations are faster than MD simulations, however, as in other coarse-graining methodology (CAC and CADD), the time step of the simulation is imposed by the events on the atomic scale domain. Despite the technical advantages of CAC over CADD, the studies have been limited to 2D or 2.5D domains. On the other hand, CADD and CAC need to reconstruct the GB structure at the atomic scale and are limited to materials with available interatomic potentials.
Another mesoscopic tool available to study GB-dislocation interactions is the phase field dislocation dynamics (PFDD) model [Koslowski et al.,, 2002, Beyerlein and Hunter,, 2016]. In this model, dislocations evolve by minimizing the total energy, hence, their interactions and reactions follow an energetically favorable pathway. Lei et al., [2013] extended the PFDD formulation to heterogeneous materials using virtual strains that describe the effect of heterogeneities and an additional evolution equation to evolve the virtual strains [Wang et al.,, 2002]. In Lei et al., [2013], the authors modeled heterogeneous materials with evolution of dislocations and voids. In Zeng et al., [2016], PFDD was used for the first time to model bimetal interfaces using PFDD with virtual strains. In Zeng et al., [2016], a misfit strain was introduced to describe the GB with a decay function that reduces the effect of the misfit strain inside the grains, while the misfit strain was derived using linear elasticity and plane stress [Hoagland et al.,, 2002], it can also be computed from molecular statics (MS) [Kharouji et al.,, 2024]. More recently, [Ma et al.,, 2022] studied an incoherent twin boundary (ITB) consisting of an array of three partial dislocations. The ITB material parameters and structure were informed by MS. GB-dislocation interaction events were in good agreement between PFDD and MD. This methodology allows the method to describe GB structures in detail and study GB-dislocation interactions precisely; however, is limited to certain systems where the GB structure can be defined in terms of well-defined dislocation arrays.
The goal of this paper is twofold, first, we present a new PFDD formulation for heterogeneous materials coupled to a micromechanical solver. The results presented here use the basic scheme introduced by Moulinec and Suquet, [1994], but can be used with other schemes. The second goal is focused on the simulation of GB-dislocations interactions using an approach that can account for various GB structures without being limited to well-defined GB structures where the dislocation arrays can be determined a priori. To achieve this, we have performed simulations of bicrystals where the GB is defined as an inclusion. The simulations parameters are informed by MS [Mishra et al.,, 2024], and compared to recent MD simulations [Dang et al.,, 2025]. In particular, we focus on the observations made by Dang et al., [2025]: GB-dislocation interaction varies depending on the applied stress, misorientation, and whether it is a minimum energy or metastable GB. MD simulations show that interactions can vary and involve only the leading partial during transmission (partial transmission). To capture this, PFDD simulations were performed using partial dislocations [Hunter et al.,, 2013]. In Section 2, the PFDD formulation is presented. This section describes how the micromechanical problem is solved, and how the total strain is used in the elastic energy terms that describe the evolution of the dislocations. Next, Section 3 presents the PFDD simulation results of dislocation reactions with different GBs in Cu. Finally, concluding remarks are given in Section 4.
2 Phase field dislocation dynamics formulation coupled with the Lippmann-Schwinger equation
In the PFDD formulation, dislocations are described using a scalar order parameter, , that is defined on each active slip system, . and correspond to a perfect crystal and a crystal that has undergone slip, while non-integer values represent a distorted crystal structure (i.e., the location of the dislocation core). The coupled formulation uses two different sets of equations, one to obtain the total strain through solving the Lippmann-Schwinger equation and the other to obtain the evolution of dislocations in the system following the pseudo-algorithm shown in Figure 1. The different components of the algorithm are described in sections below.
2.1 Dislocation mechanics solving the Lippmann-Schwinger equation
The stress of the heterogeneous domain at a point is computed as
where is the fourth order stiffness tensor of the phase or material at point and is the elastic strain. The elastic strain is decomposed as
where the total strain is split into a fluctuation term and a prescribed volume average term corresponding to the homogeneous strain, and is the plastic strain. The plastic strain depends on the order parameter as [Koslowski et al.,, 2002, Beyerlein and Hunter,, 2016]
where is the Burgers vector magnitude, is the interplanar spacing, is the number of active slip systems, and, and are the slip direction and slip plane normal of slip system . Introducing a reference material with stiffness and using the decomposition of the elastic strain, the stress may be redefined as
The last two terms in the right hand side of Equation 4 form the polarization tensor, defined as
or more generally
Using the polarization stress and Equation 4, the balance of linear momentum gives the following heterogeneous partial differential equation
Introducing the Green’s function, the above equation can be solved in the form of an integral Lippmann-Schwinger equation for the unknown strain:
where denotes spatial convolution and is the modified Green tensor defined as the second derivative of the Green’s function . In this work, Equation 8 is solved using the basic scheme proposed by Moulinec and Suquet, [1994] and reference stiffness tensor , where and are the minimum and maximum elastic properties, see Algorithm 1 in the Supplementary Material.
2.2 Dislocation evolution using the time-dependent Ginzburg-Landau (TDGL) kinetic equation
The total energy is composed of three contributions [Koslowski et al.,, 2002, Beyerlein and Hunter,, 2016]:
where , and are the elastic, lattice and external energy. Note that in some PFDD formulations there is an additional gradient energy term representing the excess energy generated by the non-uniformity of the dislocation [Xu et al.,, 2019, Shen and Wang,, 2003, Kim et al.,, 2021]. The elastic energy accounts for the elastic interactions between dislocations and is defined as
To define the elastic energy of the heterogeneous system, the elastic strain can be derived from Equation 4 using Green’s functions and the decomposition of the total strain:
and in Fourier space
where is the wave number vector and the total strain is obtained by solving the Lippmann-Schwinger equation, see Equation 8. By applying the Plancherel theorem and using the definition of the elastic strain in Fourier space, the elastic energy is defined as
where the interaction matrices , and are defined as
The lattice energy represents the energy required to move the dislocation core through the crystal lattice by the breaking and reforming of atomic bonds. It can be expressed using different forms such as piecewise quadratic functions [Koslowski et al.,, 2002], sinusoidal functions [Xu et al.,, 2020, Peng et al.,, 2020] or Fourier series [Hunter et al.,, 2011, Schoeck,, 2001, Shen and Wang,, 2004], depending on the simulation parameters:
where are the periodic potentials. For simulations in this work, it is necessary to use partial dislocations. For this, we consider the slip plane and three different slip directions with order parameters, , and , that correspond to the directions , and , respectively [Hunter et al.,, 2013].
The external energy accounts for interactions between the applied stress and dislocations and is defined as
where is the applied stress. The homogeneous strain is determined by the boundary conditions. For a stress-free system, the homogeneous strain is determined by minimizing the total strain energy with respect to the homogeneous strain
Hence, the homogeneous strain is
where , , , and . In a homogeneous system, Equation 20 is redefined as which simplifies to the definition of the external energy defined in the homogeneous PFDD formulation [Jin and Khachaturyan,, 2001].
The system evolves using a time-dependent Ginzburg-Landau (TDGL) kinetic equation, which minimizes the total energy
where is a coefficient that represents the mobility of the dislocations.
The lattice energy in Equation 17 is redefined as
where is implemented as a Fourier series [Hunter et al.,, 2011, Schoeck,, 2001, Shen and Wang,, 2004]:
The expansion coefficients and can be obtained by fitting the -surface computed using MD or DFT, in this work we used the same coefficients computed by Ma et al., [2022].
2.3 Representation of grain boundaries
GBs are represented as inclusions with different properties compared to the surrounding grains. We use MS calculations to define three GB properties: (1) width (2) second-order elastic constants (SOECs), and (3) lattice energy. The width of the GB was determined using a two-class Gaussian Mixture Model clustering of the atomistic configuration as explained by Mishra et al., [2024]. To determine the SOECs of the GB region, we performed affine deformation of a bicrystal with 3D periodic boundary conditions. A prescribed set of lattice strains () was applied to an energy minimized bi-crystal under the assumption of general elastic anisotropy. Per-atom energies for the atoms in the GB region were used to calculate the energy density in the GB region as a function of the applied strain. The elastic stiffness coefficients were determined from these energy density-strain data. We assume that the strain in the bi-crystal is homogeneous and equal to the applied strain. The Lamé constants are determined from the SOECs using the Voigt average (See Supplementary Figure 1).
We perturb the lattice energy in the GB region using an additional term that accounts for the difference in atomic structure and its effect on the evolution of the dislocation core. This additional energy barrier is relevant to the resulting dislocation-GB interaction; however, it cannot be determined analytically but it can be estimated from atomic scale simulations. Previous studies have estimated the energy barrier using nudged elastic band (NEB) method [Tsuru et al.,, 2009] and MD simulations [Sangid et al.,, 2011, Li et al.,, 2020, 2022]. In Sangid et al., [2011] and Li et al., [2020], the energy barrier and GB energy are correlated using a power law equation; however, the number of data points used are not enough to create a robust model. Thus, in this work, we take a different approach and use the GB energy to scale the lattice energy in the bulk grains (see Equation 22) to define the perturbation on the lattice energy in the GB region as
where is a fitting constant, is the GB energy, is the unstable stacking fault energy, is the shear modulus of the grain and is the shear modulus of the GB. This term acts as a perturbation of the -surface in the bulk grain to account for the disordered nature of the GB with respect to the bulk crystal so the total energy at the GB region is calculated as . To align with previous work, this term could be modified as where and are fitting parameters [Sangid et al.,, 2011, Li et al.,, 2020]. We highlight again that the energy barrier to dislocation glide within a GB due to the atomic disorder is still an open research question, and is expected to have strong dependence on the GB structure. Thus, these equations are representative of this effect, but may need to be redefined to GB structures not considered in this paper. In particular, specialized GB structures such as cube-on-cube orientation or coherent twin boundaries with low values of GB energy , are likely examples of cases in which different descriptions of this energy barrier may be necessary.
2.4 Material properties and simulation setup
In previous work [Dang et al.,, 2025], dislocation-GB interactions were studied with MD for several different symmetric tilt GBs in Cu including both stable and metastable GB configurations. From this database of GB structures [Dang et al.,, 2025], we selected two 110 symmetric tilt GBs and their respective metastable states to consider in this paper. The first GB is a low angle GB (LAGB) with (1,1,6) GB plane and a misorientation angle of 26.5∘ and the second GB is a high angle GB (HAGB) with (5,5,2) GB plane and a misorientation angle of 148.4∘. These GBs have different GB widths, GB energies, elastic constants and GB-dislocation reactions as predicted with MD, see Table 1. The Lamé constants of the Cu grains surrounding the GB region are and and mJ/m2. For each GB, the minimum energy structure plus two higher metastable structures were selected for this study. We note that we were unable to determine the SOECs for some metastable GBs as they were unstable under some deformation modes. Therefore, the Lamé constants and GB width of the minimum energy GB structure is used also for the metastable cases.
| Type | GB Energy | Width | MD Reaction | |||
| mJ/m2 | GPa | Å | 250 MPa | 500 MPa | 750 MPa | |
| LAGB | 751 | 29.16, 119.78 | 18.69 | A/P | T | T |
| LAGB | 1003 | 29.16, 119.78 | 18.69 | A/P | T | T |
| LAGB | 1228 | 29.16, 119.78 | 18.69 | A/P | A/P | T |
| HAGB | 847 | 40.11, 112.11 | 12.6 | A/P | A/P | T |
| HAGB | 938 | 40.11, 112.11 | 12.6 | A/P | A/P | T |
| HAGB | 998 | 40.11, 112.11 | 12.6 | A/P | A/P | A/P |
A domain of 256x4x128 is used with a grid size that corresponds to the Burgers vector norm of a screw dislocation in Cu ( Å). Periodic boundary conditions are used in the three directions and the domain size is chosen such that the periodic interactions are minimized. An inclusion, representing the GB region, is located in the center of the domain and the position of the initial dislocation dipole is presented in Figure 2. A screw dislocation dipole is placed in the left grain, with the left monopole placed at 63 grid points from the domain boundary with a distance of 50 grid points between the dislocations. The domain is subjected to three different applied shear stresses MPa, resulting in a different reaction for a given GB energy value as shown in Table 1. The dislocation mobility is set to with .
3 Results and discussions
The reactions in Table 1 are reproduced using the PFDD model by setting for each type of GB. The values used in the following simulations for the LAGB and HAGB are 0.04 and 0.086, respectively. As the elastic constants and energies are fixed for a given GB, Equation 24 may be redefined as
with
The value of is shown in Table 2 for each GB. Based on the reactions at different applied stress, the six GBs can be divided into three categories as shown in Table 2. One can notice the value of is close within each category. If is between 0.213 and 0.304, the GB-dislocation reactions are (A/P, T, T), if is between 0.361 and 0.41, the GB-dislocation reactions are (A/P, A/P, T), and if is equal or higher than 0.442, the GB-dislocation reactions are (A/P, A/P, A/P). The values roughly correlate with strain-gradient metrics from the strain functional descriptors used in the two-class Gaussian Mixture Model clustering in [Mishra et al.,, 2024]. More details are provided in the supplemental information.
| Type | GB Energy | PFDD Reaction | ||||
| mJ/m2 | 250 MPa | 500 MPa | 750 MPa | |||
| LAGB | 751 | 0.04 | 0.213 | A/P | T | T |
| LAGB | 1003 | 0.04 | 0.304 | A/P | T (C1) | T |
| LAGB | 1228 | 0.04 | 0.386 | A/P | A/P (C2) | T |
| HAGB | 847 | 0.086 | 0.361 | A/P | A/P (C3) | T (C4) |
| HAGB | 938 | 0.086 | 0.410 | A/P | A/P | T |
| HAGB | 998 | 0.086 | 0.442 | A/P | A/P | A/P |
Among the cases shown in Table 1, four different cases are analyzed here in more detail: (C1) LAGB with GB energy mJ/m2 under an applied stress of MPa, (C2) LAGB with GB energy mJ/m2 under an applied stress of MPa, (C3) HAGB with GB energy mJ/m2 under an applied stress of MPa and (C4) HAGB with GB energy mJ/m2 under an applied stress of MPa. Note that the other cases have been simulated and the GB-dislocation reactions from MD are reproduced, see Supplementary Figures 2-15. C1, C2 and C3 have different energies and same applied stress while C3 and C4 are the same GB with different applied stresses. for C1, C2 and C3/C4 is 0.304, 0.386 and 0.361, respectively (see Table 2). To the authors knowledge, there is no way to compute analytically the values of , for the work presented here the values are chosen to reproduce the MD reactions.
Figures 3, 4, 5 and 6 show the change of the disregistry and stress evolution at different time steps for cases C1, C2, C3 and C4. The disregistry is computed as [Shen and Wang,, 2004]
where is the Burgers vector direction of the initial perfect dislocation. The dislocation transmits in cases C1 and C4, and it is pinned in cases C2 and C3 as the dislocation stays blocked inside (Figure 4) or next (Figure 5) to the GB region. Three main observations need to be highlighted, (i) transmission is easier for LAGB which is in accordance with MD simulations and experimental observations [Nieto-Valeiras et al.,, 2024, Adams et al.,, 2019], (ii) the change in GB structure between C1 and C2 is reflected in the increase of GB energy and this additional GB energy results in an increased resistance to the dislocation and results in the dislocation being pinned, and (iii) the change in the interaction between cases C3 and C4 from being pinned to transmitted is a direct effect of the applied stress. Also, in C4 the dislocation evolves faster due to the higher applied stress of 750 MPa.
The dislocation disregistry change during the interaction is shown in Figure 7. Moreover, the change in the stacking fault width is more clear by plotting its derivative along the slip direction, see Figure 8. These detailed views of the disregistry are shown at the same time steps in Figures 3, 4, 5 and 6. The initial value of the stacking fault width in the four cases is 4 or 5 grid points that correspond to 1nm which is in agreement with previous studies [Jin et al.,, 2006]. For the cases where the dislocation transmits across the GB, there is a difference in the stacking fault width during the interaction. In case C1, the stacking fault width remains constant with values between 4 and 5 grid points, see Figures 3, 7a and 8. While in case C4, the stacking fault width decreases to 2 grid points (Figure 6b) and comes back to its initial shape (Figure 6f). Something similar happens in the cases that get pinned. In case C2, the dislocation gets pinned with the trailing partial being blocked at the interface between the left grain and the GB region and the leading partial stays inside the GB region, also the stacking fault width changes slightly to 3 grid points as shown in Figures 7b and 8. In case C3, as the dislocation approaches the GB and gets pinned the stacking fault width decreases to half of its original width from 4 to 2 grid points, see Figures 7c and 8. Due to this difference, the interaction in case C2 may be categorized as absorption instead of pinning. Note that absorption is a challenging interaction to model as the interaction can have different scenarios, i.e., dislocations can glide and re-emit later, or the dislocation can change the GB structure leading to nucleation of new dislocations [Jin et al.,, 2006, Dang et al.,, 2024, Liang et al.,, 2019, Wang,, 2015]. However, modeling absorption and subsequent dislocation glide along the GB plane is not available in the present PFDD model but could be considered by adding existing glide and cross slip PFDD extensions [Fey et al.,, 2022, Wise et al.,, 2025].
In summary, the variation in the stacking fault width is higher in C3 and C4 as shown in Figures 7 and 8. As is added to the total energy, it acts as an additional attractive force between the two partial dislocations which explains the change in the disregistry. Hence, the difference between the stacking fault width in C1, C2, C3 and C4 is proportional to despite of the higher applied stress in C4. In the Supplementary Material, is plotted against averaged strain gradient metrics showing good correlation with both, net strain gradient and gradient on uniaxial strain metrics computed using MS [Mishra et al.,, 2024]. Similar changes in the stacking fault width have been reported in MD simulations by Jin et al., [2006] (cf. Figure 2) where the partial dislocation is constricted to an almost perfect dislocation at the GB before transmission or absorption of the partial dislocation, and Dang et al., [2024] (cf. Figure 2) where the partial dislocation stacking fault width reduces to almost half of its original width when the dislocation is pinned.
4 Conclusion
A new PFDD formulation that couples phase field and micromechanical solvers is presented. This new formulation allows for the study of heterogeneous materials with different stiffness tensor by solving the Lippmann-Schwinger equation. The total strain obtained from the Lippmann-Schwinger equation is one of the terms in the elastic energy density used within the time-dependent Ginzburg-Landau (TDGL) kinetic equation.
In this work, GB-dislocation interactions are studied by representing GBs as inclusions with an specific width obtained from MS simulations [Mishra et al.,, 2024]. Two additional parameters are added to describe GBs, (i) an elastic energy which is considered by computing a different GB elastic stiffness tensor and (ii) a lattice energy that represents the effect of the disordered atomic structure within the GB region. Two different groups of GBs are studied including stable and metastable states: LAGB with GB plane (1,1,6) and misorientation angle of 27∘ and HAGB with GB plane (5,5,2) and misorientation angle of 178∘. The GB-dislocations interactions predicted using PFDD show excellent agreement with the MD results. However, these results depend on the calibration parameter which varies as a function of the misorientation. We presented an analysis on the stacking fault width, the initial width and reduction of the width during the interaction which shows good agreement with MD simulations [Jin et al.,, 2006, Dang et al.,, 2024]. Finally, by adding an additional lattice energy for the GB region proportional to , we can account for the dependency of GB-dislocation interaction on the GB energy or GB structure.
In contrast to the sharp interface method [Zeng et al.,, 2016, Ma et al.,, 2022], this new methodology opens the door to model more general GB-dislocation interactions. Future work will be focused on the improvement of the GB description and the inclusion of other interactions such as GB absorption and glide of the dislocation across the GB. Also, other GB-dislocation interactions such as reflection should be considered.
Acknowledgments
AH, BM, and NM acknowledge support from the Physics and Engineering Models (PEM) Materials project within the Advanced Simulation and Computing (ASC) Program at LANL (ASC-PEM-Materials Project). AM and NM also acknowledges funding from the LDRD-DR “Investigating How Material’s Interfaces and Dislocations Affect Strength” (iMIDAS) under Grant No. 20210036DR (Abigail Hunter, PI and Saryu J. Fensin, Co-PI) and for the initial MD calculations and methodology. BM also acknowledges Ian Chesser and Ian Wise for their constructive comments and discussions during the development of the work. LA-UR-25-26159.
Supplementary material
Basic scheme
The fourth-order Green operator is given by
In the particular case of isotropic elasticity it may redefined as
Details of the coupled formulation
This sections aims to show some of the steps that were not shown in Section 2 to get Equations 11, 13 and 20.
Equation 11 is obtained by following these steps: From Equation 4 and applying the balance of linear momentum one gets:
rearranging the equation above and getting the total strain multiplied by the reference elastic modulus on the left hand side, one gets
using Green’s operator and the Fourier transform the solution for the total strain is
Equation 11 is obtained by replacing the equation of total strain presented above into the strain decomposition (Equation 2):
The Fourier transform of the elastic strain is used to get the equation for the elastic energy
replacing the elastic strain into the equation above, the elastic energy may be redefined as
multiplying the two elastic strains the elastic energy gives the same expression shown in Equation 13
Finally, Equation 20 is obtained through the derivation of the strain energy with respect to the average strain. The strain energy is defined as:
making the derivative of the strain energy with respect to the average strain equal to zero and using the major symmetry of the elastic modulus:
rearranging the terms and dividing by the volume
note that by definition the term . Equaton 20 is obtained using the definitions of averaged quantities presented by the end of Section 2 (, , , and ). For the case of applied strain and . If , the average strain is defined as
which is equivalent to the formulation for a homogeneous material.
Calculation of second order elastic constants
Phase field dislocation dynamics simulations
In the main text only three cases were presented, the remaining simulations are shown in the Supplementary Figures 2-15.
Averaged strain-gradient metrics calculated for classification in [Mishra et al.,, 2024] is plotted as a function of in Figures 16 and 17. These were computed using the small atomistic simulation cells used for GB width and elastic stiffness calculations. The average values were computed over a region around the GB corresponding to the GB width. These show that the are correlated with strain-gradient metrics. These GBs had the same macroscopic degrees of freedom and similar energy values to the GBs used in dislocation-GB interaction study Dang et al., [2025].
The data is more noisy when computed from large simulation cells as shown in Figures 18 and 19. In this case the averages were computed using all atoms in a region around the dislocation interaction site, defined by the GB width (normal to the GB plane) and 60 Å(normal to the tilt axis). As the larger atomistic simulation cells are less constrained, the GB undergoes relaxation resulting in differences in energy values compared to the small cells.